OpenStudy (anonymous):

Evaluate the definite integral of (x-5)^21 dx from 0 to 10

hartnn (hartnn):

put x-5 = u du=... ?

OpenStudy (anonymous):

1? because isn't du just derivative of x-5?

hartnn (hartnn):

not exactly 1 when x-5 = u 1-0 = du/dx so, du = dx got this ?

OpenStudy (anonymous):

Yes and then would it be (u)^21 du?

hartnn (hartnn):

yes! now don't forget to change the limits when x= 0, u=... ? when x = 10, u =.... ?

OpenStudy (anonymous):

-5? and 5?

hartnn (hartnn):

yes! so, you have \(\int \limits_{-5}^5u^{21}du\) tight ? any ideas how to solve this ?

OpenStudy (anonymous):

could you put the 5 and -5 into x-5? and take antiderivative of u^21? and put 5 and -5 in and subtract?

hartnn (hartnn):

can you think of an easier way ? i mean look at the limits -5 to +5 sounds like -a to +a do you recollect any property of definite integral ??

OpenStudy (anonymous):

i dont remember one that had a and -a in it

hartnn (hartnn):

ok, let me tell ya, \(\huge \int \limits_{-a}^a f(x)dx = 0\) IF f(x) is odd function. so, here IS u^21 an ODD function ? do you know what is an odd function ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

yes for what ?

OpenStudy (anonymous):

is it odd because you can get f(x)=-f(x)?

hartnn (hartnn):

yes, its odd, so it makes the integral = 0 simple, right ?

OpenStudy (anonymous):

so it's just a basic property?

hartnn (hartnn):

yup.

OpenStudy (anonymous):

that's much simpler than I had thought.

OpenStudy (anonymous):

Thank you for explaining this to me!

hartnn (hartnn):

welcome ^_^